Need all parts some hints are giving ! If you don’t know don’t answer need experts only
Question # 7 A random variable x has the following distribution: px) 0. -2 -1 0 +1 +2 +3 0.2 k 0.3 2k (i) Find k. (ii) Find P(x <0). (iii) Find P Answer-hints: (iii) P(-2 < x < +2) = 0.8; (v) E11-1.01 2 x +2). (iv) Plot CDF of x; and (v) Find E[x]
Answer #1

i)

for sum of probabilties should sum up to 1

hence 0.1+k+0.2+k+0.3+2k=1

4k=0.4

k=0.1

ii)P(X<0)=P(X=-2)+P(X=-1)=0.1+0.1=0.2

iii)P(-2<X<2)=P(X=-2)+P(X=-1)+P(X=0)+P(X=1)+P(X=2)=0.1+0.1+0.2+0.1+0.3=0.8

iv)

0.8 F(x) 0.6 04_ 4

v)

E(X)=sumxP(x)=-2*0.1+(-1)*0.1+0*0.2+1*0.1+2*0.3+3*0.2=1